OpenStudy (anonymous):

If the width of a rectangular piece of wood is x + 2 inches long and the length is three times as long with an additional inch added, what is the area of the wood? A. 3x² + 12x + 12 B. 3x² + 12x + 14 C. 3x² – 13x + 14 D. 3x² + 13x + 14

OpenStudy (anonymous):

Nvm, it was D. Thanks anyways @Jim766

OpenStudy (jim766):

A = L * W W = x + 2 L = 3(x+ 2) + 1= 3x + 6+1 = 3x +7 A = L * W A = (3x + 7)(x + 2) now foil it